OpenStudy (anonymous):

Write an equation in slope intercept form of the line that passes through (4, 1) and (5, -1). A. y = -4x + 1 B. y = -1/2 x + 9 C. y = 4x + 1 D. y = -2x + 9

OpenStudy (shamil98):

Use the slope formula: \[\large \frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (shamil98):

Then input your values into the point-slope form and solve for y ( to get your slope-intercept form). \[\large y-y_1 = m(x-x_1)\]

OpenStudy (anonymous):

@shamil98 I did that and got y= 0x +1.. that's not an answer so....

OpenStudy (the_fizicx99):

What is -1 - 1 ?

OpenStudy (the_fizicx99):

@taylorwristshot

OpenStudy (anonymous):

@tHe_FiZiCx99 -2

OpenStudy (the_fizicx99):

thats your m (slope) so now plug in your value's for y-y1=m(x-x1)

OpenStudy (anonymous):

@tHe_FiZiCx99 So that would be 4-1= -2 (5- -1) right?

OpenStudy (the_fizicx99):

No, Y - Y1=m(X-X1 Just plug in for: Y1, M and X1

OpenStudy (anonymous):

@tHe_FiZiCx99 Oh okay.. then it would be Y-4= -2 (x-5)

OpenStudy (the_fizicx99):

No, ordered pairs are (x,y) try again

OpenStudy (anonymous):

@tHe_FiZiCx99 Im so confused now...

OpenStudy (the_fizicx99):

Ok, ordered pairs are shown as (x,y) you have (4,1) where 4 is your x and y is your 1

OpenStudy (the_fizicx99):

Then you have (5,-1) again, where x is the 5 and -1 is the y

OpenStudy (the_fizicx99):

Understand?

OpenStudy (the_fizicx99):

So again, what is : y-y1=m(x-x1)

OpenStudy (the_fizicx99):

@taylorwristshot

OpenStudy (anonymous):

@tHe_FiZiCx99 I guess.. Then wouldn't it be 1-y1= -2 (4-x1)?

OpenStudy (the_fizicx99):

No, the formula is : y-1=-2(x-4)

OpenStudy (the_fizicx99):

I only plugged in the y1 and x1

OpenStudy (anonymous):

@tHe_FiZiCx99 Oh.. alright.. but that's not even close to being an answer...

OpenStudy (the_fizicx99):

Solve it,

OpenStudy (the_fizicx99):

So what'd you get?

OpenStudy (the_fizicx99):

I'm gonna assume you solved it :) y-1=-2(x-4) y-1=-2x - 8 +1 + 1 y = -2x + 9

OpenStudy (anonymous):

@tHe_FiZiCx99 YES! That's what I got. Sorry for the late reply, was busy. Thank you!!!

OpenStudy (the_fizicx99):

Yw.